The left and right page numbers of an open book are two consecutive integers whose sum is 281. find these page numbers
The smaller page number is?

the larger page number is?

Answer by faceoff57(108)

The left and right page numbers of an open book are two consecutive integers whose sum is 281. find these page numbers.
X + (x + 1) = 281
X + x + 1 = 281
2x + 1 = 281
2x = 281 - 1
2x = 280
X = 140
The smaller page number is? 140
The larger page number is? 141